OpenStudy (anonymous):

A box of crayons costs $1.75, including tax. Mr. Valentino wants to purchase boxes of crayons for his class and has a $25 budget. Write an inequality to solve for the number of boxes of crayon Mr. Valentino can purchase with his budget.

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

Mr Valentino will spends less than or equal to the amount of money he has

OpenStudy (anonymous):

$1.75x ≤ $25 $1.75x ≥ $25 $25x ≤ $1.75 $25x ≥ $1.75

OpenStudy (anonymous):

these r my options

OpenStudy (unklerhaukus):

so I guess x represents the number of boxes of crayons

OpenStudy (unklerhaukus):

the total spent is the cost per box times the number of boxes

OpenStudy (anonymous):

so whats the answer? im very confused

OpenStudy (unklerhaukus):

say for instance Mr Valentino was to purchase 1 box, 1 would be $1.75 \(\times\) 1 = $1.75 but Mr Valentino want to get as more boxes 2 boxes would be $1.75 \(\times\) 2 = $3.50 . . . x boxes would be $1.75 \(\times\) . . .

OpenStudy (anonymous):

$1.75x ≥ $25????

OpenStudy (unklerhaukus):

Mr Valentino will spends less than or equal to the amount of money he has the cost is on the left hand side, the money he has is on the right

OpenStudy (anonymous):

so i got it right then? i think so

OpenStudy (unklerhaukus):

not quite,

OpenStudy (anonymous):

here: $1.75x ≤ $25

OpenStudy (unklerhaukus):

yer does it make sense now?

OpenStudy (unklerhaukus):

Now you have to solve the inequality,

OpenStudy (anonymous):

Solve for x: −2x + 5 < 7 x > −1 x < −1 x > −6 x < −6

OpenStudy (unklerhaukus):

is that you on biochat?

OpenStudy (anonymous):

hey, i really uck at this

OpenStudy (anonymous):

oh god

OpenStudy (unklerhaukus):

we haven't finished the first question, divide both sides of the inequality by $1.75 to get x≤ ... this is the the number of boxes of crayons Mr Valentino can purchase

OpenStudy (anonymous):

no no no, i dont need to solve them